OpenStudy (anonymous):

Plug x = 2 into this equation to solve for y. y = 5x + 6 A. y = 16 B. y = 4 C. y = 11 D. y = 10

OpenStudy (whpalmer4):

Do as it says: \[y = 5x + 6\]\[y = 5(2) + 6\]\[y = 10+6\]\[y = \]